OpenStudy (anonymous):

please hellp me A kilobyte is 210 bytes and a megabyte is 220 bytes. How many kilobytes are in a megabyte? 2200 kilobytes 230 kilobytes 210 kilobytes 2 kilobytes

OpenStudy (anonymous):

Umm, this question doesn't really make sense. I get an answer of 1.04 kilobytes. Did you type it in right?

OpenStudy (anonymous):

yeah i dont get it eather its typed exact

OpenStudy (anonymous):

@Mcwoodman

OpenStudy (anonymous):

Okay, I'm not sure. The closest answer is 2 kilobytes so I would expect that to be the answer.

OpenStudy (anonymous):

i think it's 2^10 for a kB, and 2^20 for a MB...

OpenStudy (anonymous):

im not sure but sounds about right my teacher hasnt explained it

OpenStudy (anonymous):

1 kB = 2^10 Bytes = 1024 Bytes 1 MB = 2^20 Bytes = (2^10)(2^10) Bytes = (1024)^2 Bytes = 1048576 Bytes

OpenStudy (anonymous):

so its gonna be 2^10 right

OpenStudy (anonymous):

@Orion1213

OpenStudy (anonymous):

yup you're right... :-)

OpenStudy (anonymous):

thank you sooo much